OpenStudy (vortish):

Complete the following parts. 2x-3y = 6 (A) write the equation in slope intercept form (B) Give the slope and y intercept of the line

OpenStudy (anonymous):

a: y=(2/3)x-2 b. slope: 2/3 , y intercept : -2

OpenStudy (vortish):

zordoloom can you show me how you got to that pls

OpenStudy (anonymous):

Don't want to step on anyone's toes, but to help... first look at ur equation and solve for y. 2x-3y=6 subtract 2x from each side. you get -3y= -2x+6 then you want to single out y, so you divide each side by -3 that makes it y=2/3x-2 the -2 divided by the -3 makes it a positive fraction

OpenStudy (anonymous):

then in the expression y=mx+b, m represents the slope and b represents the y intercept

OpenStudy (vortish):

thanks rosey